$SPAD/src/input nqip.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool nqip.output
)set message test on
)set message auto off
)clear all

———-

Note that the conversions of results from DoubleFloat to Float will become unnecessary if outputGeneral is extended to apply to DoubleFloat quantities.

— * —

--S 1 of 14
outputGeneral 5
--R
--R                                                                   Type: Void
--E 1
--S 2 of 14
xvals := [0.00,0.04,0.08,0.12,0.22,0.26,0.30,0.38,0.39,0.42,0.45,
               0.46,0.60,0.68,0.72,0.73,0.83,0.85,0.88,0.90,1.00];
--R
--R
--R                                                            Type: List(Float)
--E 2
--S 3 of 14
yvals := [4.0000,3.9936,3.9746,3.9432,3.8135,3.7467,3.6697,3.4943,
                 3.4719,3.4002,3.3264,3.3017,2.9412,2.7352,2.6344,
                        2.6094,2.3684,2.3222,2.2543,2.2099,2.0000];
--R
--R
--R                                                            Type: List(Float)
--E 3
--S 4 of 14
result := nagPolygonIntegrate(xvals,yvals);
--R
--R   There are no library operations named nagPolygonIntegrate
--R      Use HyperDoc Browse or issue

--R                        )what op nagPolygonIntegrate
--R      to learn if there is any operation containing "
--R      nagPolygonIntegrate " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagPolygonIntegrate with argument type(s)
--R                                 List(Float)
--R                                 List(Float)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 4
--S 5 of 14 used to work?
result.integral :: Float
--R
--R   There are no exposed library operations named result but there is
--R      one unexposed operation with that name. Use HyperDoc Browse or
--R      issue
--R                             )display op result
--R      to learn more about the available operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      result with argument type(s)
--R                             Variable(integral)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 5
--       3.1414
--S 6 of 14 used to work?
result.errorEstimate :: Float
--R
--R   There are no exposed library operations named result but there is
--R      one unexposed operation with that name. Use HyperDoc Browse or
--R      issue
--R                             )display op result
--R      to learn more about the available operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      result with argument type(s)
--R                           Variable(errorEstimate)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 6

--       - 0.000025627
--S 7 of 14
coords := transpose matrix [xvals, yvals];
--R
--R
--R                                                          Type: Matrix(Float)
--E 7
--S 8 of 14
result := nagPolygonIntegrate coords;
--R
--R   There are no library operations named nagPolygonIntegrate
--R      Use HyperDoc Browse or issue
--R                        )what op nagPolygonIntegrate
--R      to learn if there is any operation containing "
--R      nagPolygonIntegrate " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagPolygonIntegrate with argument type(s)
--R                                Matrix(Float)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 8
--S 9 of 14 used to work?
result.integral :: Float
--R
--R   There are no exposed library operations named result but there is
--R      one unexposed operation with that name. Use HyperDoc Browse or
--R      issue
--R                             )display op result
--R      to learn more about the available operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      result with argument type(s)
--R                             Variable(integral)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 9
--       3.1414
--S 10 of 14 used to work?
result.errorEstimate :: Float
--R
--R   There are no exposed library operations named result but there is

--R      one unexposed operation with that name. Use HyperDoc Browse or
--R      issue
--R                             )display op result
--R      to learn more about the available operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      result with argument type(s)
--R                           Variable(errorEstimate)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 10
--       - 0.000025627
--S 11 of 14 broken
nagPolygonIntegrate([1,2,3],[1,2,3,4])
--R
--R   There are no library operations named nagPolygonIntegrate
--R      Use HyperDoc Browse or issue
--R                        )what op nagPolygonIntegrate
--R      to learn if there is any operation containing "
--R      nagPolygonIntegrate " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagPolygonIntegrate with argument type(s)
--R                            List(PositiveInteger)
--R                            List(PositiveInteger)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 11
--
--   Error signalled from user code:
--      The lists supplied to nagPolygonIntegrate are of different
--      lengths: 3 and 4.
--S 12 of 14 broken
nagPolygonIntegrate([[1,2,3],[4,5,6]])
--R
--R
--R   (5)  nagPolygonIntegrate
--R                           [1,2,3],[4,5,6]
--R                                                                 Type: Symbol
--E 12
--
--   Error signalled from user code:
--      Please supply the coordinate matrix in nagPolygonIntegrate with
--      each row consisting of single a x-y pair.

--S 13 of 14
outputGeneral()
--R
--R                                                                   Type: Void
--E 13
--S 14 of 14
output "End of tests"
--R
--R   End of tests
--R                                                                   Type: Void
--E 14
)spool
)lisp (bye)

———-

References^{[1] nothing}

8